OpenStudy (anonymous):

Give the equivalent numerator if the denominator is (a - 6)(a - 1)(a + 6). 6a² - 42a + 36 6a - 6 6a + 36

OpenStudy (anonymous):

\[\frac{ 6 }{ a^2-7a+6 }\]

OpenStudy (anonymous):

@Owlcoffee

OpenStudy (owlcoffee):

well, here it is a little trickier, we have to factorize the denominator first.

OpenStudy (anonymous):

Okay..lol

OpenStudy (owlcoffee):

So, in order to factorize, you'll have to make the denominator equal zero and use the quadratic formula to find the root, but I won't do that, I'll leave it to you as a Homework. So, facotrized it looks like this: \[\frac{ 6 }{ (a-6)(a-1) }\] And once again, multiply by the expression we require: \[\frac{ 6(a+6) }{ (a-6)(a-1)(a+6) }\] And, distributive above: \[\frac{ 6a+36 }{ (a-6)(a-1)(a+6) }\] The answer is the numerator of this last fraction.